When water is "broken" into ions, two basic ions are formed: the hydrogen ion (H+) and the hydroxide ion (OH-). (H₂O ⇒ H⁺ + OH⁻) The hydrogen ion is the acidic part and the hydroxide ion is the basic part. If a chemist knows the pH of a solution the concentration of that solution is calculated with the equation pH = -log[H+]. A pH = 7 is a neutral solution, a pH < 7 indicates an acidic solution and a pH > 7 indicates a basic solution. The square brackets around the symbol H+ is used in chemistry to mean "the concentration of hydrogen ions".

(a) The pH of tomatoes is 4. What is the hydrogen-ion concentration?
(b) The hydrogen-ion concentration of eggs is 3.16 x 10^-8 moles/litre. What is the pH of eggs?
(c) The product of the concentration of the hydrogen ion and of the hydroxide ion, [OH-], is 10^-14. Which is expressed in the equation [H+][OH-] = 10^-14. Calculate the pH of a solution with [OH-] = 10^-5
(d) Calculate the [OH-] of a solution with a pH of 2.

Answer by josgarithmetic(39838)

-------------------------------------------------------
(a) The pH of tomatoes is 4. What is the hydrogen-ion concentration?
(b) The hydrogen-ion concentration of eggs is 3.16 x 10^-8 moles/litre. What is the pH of eggs?
(c) The product of the concentration of the hydrogen ion and of the hydroxide ion, [OH-], is 10^-14. Which is expressed in the equation [H+][OH-] = 10^-14. Calculate the pH of a solution with [OH-] = 10^-5
(d) Calculate the [OH-] of a solution with a pH of 2.
--------------------------------------------------------


I will give help for part (c) only.

You understood [H+][OH-]=10^(-14).
You are given a hydroxide concentration value and asked for pH.

[H+]*10^(-5)=10^(-14)
[H+]=10^(-14+5)
[H+]=10^(-9)

pH=9